June 25, 2018. There is a power outage scheduled at the ATDC tomorrow night to make repairs due to the flood. Pinaki will begin the Titan shutdown tomorrow morning and the scope will be back online by Wednesday afternoon. Sorry, we are as surprised as you are!
Currently browsing tag
May 8, 2018. The FESEM has a cooling system that does not depend on the process water circuit that is going off later this week. Therefore, I will be leaving the FESEM on over the 2 days that the process water will be off. SO, the FESEM will be on all …
May 4, 2018. The M&M power will not be affected by the power outage next week, but the cooling water for equipment will go off when the Dow power is shut down. Some ACMAL equipment requires that cooling water, so I will shut down the SEM, FESEM, TEM and FIB at noon on Wednesday, …
November 15, 2017. We have had cooling water problems for a few weeks. Since the last one a week ago a few parts aren’t working right. That isn’t a problem because we aren’t using that yet anyway. The scope is ready to be used in conventional TEM mode.
January 30, 2017. The ESEM is back online and ready to use. Best as we can tell right now, the EDS is working too. We will have the EDS manufacturer come in to check the system but we might as well use it in the meantime.
December 21, 2016. A broken pipe poured water onto the ESEM and computers Monday. Obviously, the microscope is down. We are going to let everything dry out over the holidays and try to restart it in the week of Jan 2.
August 12, 2016. All, I have blocked out time on 3 scopes on Wednesday next week (Aug 17) for a scheduled cooling water outage. Plan your work accordingly. Thanks, Owen
December 3, 2013. The water pump in the Denton chiller failed. I have a new on order that will be here Thursday. Provided the pump arrives Thursday, I can have the Denton back online Friday morning. If you need carbon coating before then we have the Meridian coater that can …
October 11, 2013. Facilities repaired the pump and the process water is back online! I’ll be turning the scopes back on in a few hours.
October 11, 2013. Yet another problem with the cooling water – now both circulating pumps are down.. The 6400, 4000 and FIB are off-line. Facilities is working on the pumps now. I’ll send you an update later this afternoon.